OpenStudy (anonymous):

Solve the system using an equivalent system. -2x-y=0 3x-y=10

OpenStudy (agent0smith):

Not sure what it means by an equivalent system...? But i'm guessing it means use eliminination If you multiply the first equation by -1 (multiply EVERYTHING by -1) then you can add the first and second equations together,

OpenStudy (anonymous):

solve for either x or y : -2x-y=0 Plug that answer into the 2nd equation. Then plug in both those answers into the equations to check solve for y: -2x-y=0 -y=2x y=-2x Substitute: 3x-y=10 3x-(-2x)=10 5x=10 x=2 Now check -2x-y=0 -2(2)-(-2(2))=0 -4+4=0 3x-y=10 3(2)-(-2(2))=10 6+4=10 10=10

OpenStudy (anonymous):

An equivalent system means that both of the equations have the same answers

OpenStudy (agent0smith):

^ i don't think that's what it means: http://mrpilarski.wordpress.com/2009/12/06/3-2-solving-systems-of-equations-with-equivalent-systems/ like i thought, it means use elimination (not substitution)